OpenStudy (anonymous):

The counter on a business copier read 18,678 at the beginning of the week and read 20, 438 at the end of the week, If the business was in operation 40 hours a week what was the average number of copies made each hour?

OpenStudy (anonymous):

\[\Large \frac{{20438 - 18678}}{{40}} = \frac{{{\rm{1760}}}}{{40}} = {\rm{44}}\]